### v4

@@ Abstract
In this paper we present a differential testing framework, DeltaNN, that allows us to assess the impact of different computational environment parameters on the performance of image recognition models during deployment, post training. DeltaNN generates different implementations of a given image recognition model for variations in environment parameters, namely, deep learning frameworks, compiler optimizations and hardware devices and analyzes differences in model performance as a result. Using DeltaNN, we conduct an empirical study of robustness analysis of three popular image recognition models using the ImageNet dataset. We report the impact in terms of misclassifications and inference time differences across different settings. In total, we observed up to 83% output label differences across deep learning frameworks, and up to 81% unexpected performance degradation in terms of inference time, when applying compiler optimizations.
@@ I Introduction
Overall, we find that conversions between DL frameworks significantly impacts output labels of the DNN models by up to 83%. We identify the source of the label-impacting error – small amounts of noise introduced in the weights by the framework conversion tools. The weight differences, although small, may be caused by floating-point rounding errors which can cause label changes when accumulated across the layers. On the other hand, we found that varying hardware accelerators and compiler optimizations do not affect model output but can lead to a non-negligible performance degradation with respect to inference time under specific scenarios. We observed up to 81% unexpected performance degradation in model inference times when applying certain compiler optimizations.
@@ VI-A1 RQ1: Label Sensitivity to DL Framework Conversions
The results are presented in Figure 6, showing the degree of dissimilarity between source and target models. As can be seen from the empty gray boxes, the conversion tool crashes in 10 out of the 36 conversions across the three DNN models, indicating that the conversion process failed. This happened due to compatibility issues between the conversion tool and a given model architecture, or the source or target DL framework. For instance, we could not convert any Keras models to TensorFlow, due to the tf2onnx tool being unable to handle some tensor element values. Additionally, we observe further 10 cases where the conversion succeed without crashing, but the target model gave considerably different labels (over 35%) from the source model. In particular, we observe a 83% discrepancy in the output labels when converting the MobileNetV2 model from Keras to PyTorch.
@@
Finally, the conversion of TF models to Keras gives varying results across models, with MobileNetV2 having 35% dissimilarity, ResNet101 having 4%, and InceptionV3 giving a model crash. This points to weaknesses in the conversion tool with certain DNN model architectures.
@@
We use the Analysis part of DeltaNN to explore in greater detail the cause of discrepancies across DNN model conversions. To illustrate the analysis in depth, we select one of the models and conversions that results in a discrepancy, InceptionV3 with TF as the source DL framework converted to target TFLite. The discrepancies observed across source-target is 4%, and in Figure 7 we show the class breakdown of the images which demonstrated differences, sorted by what proportion of that class showed discrepancies. We highlight a subset of the class labels on the x-axis. We observe that some classes are impacted more than others, with some classes such as “walker hound” disagreeing on 100% of the images. However, this graph is not indicating test set accuracy, instead it is about agreement between the source and converted models, which under ideal circumstances we would expect to have equal agreement in all cases.
@@ VIII Conclusion
We introduced the DeltaNN Differential testing framework to explore the impact of computational environment parameters on image recognition models. In particular, we study the effect of converting between popular deep learning frameworks (TensorFlow, Keras, TFLite, PyTorch), compiler optimization settings, and hardware devices on the output labels of three widely used image recognition models. We also monitor the impact of these parameters on model inference time. Overall, we find that conversions between DL frameworks significantly impact output labels of the DNN models by up to 83%. Our framework also provides analysis capabilities for label discrepancies stemming from framework conversions.

### v5

@@ Abstract
In this paper we present a differential testing framework, DeltaNN, that allows us to assess the impact of different computational environment parameters on the performance of image recognition models during deployment, post training. DeltaNN generates different implementations of a given image recognition model for variations in environment parameters, namely, deep learning frameworks, compiler optimizations and hardware devices and analyzes differences in model performance as a result. Using DeltaNN, we conduct an empirical study of robustness analysis of three popular image recognition models using the ImageNet dataset. We report the impact in terms of misclassifications and inference time differences across different settings. In total, we observed up to 100% output label differences across deep learning frameworks, and up to 81% unexpected performance degradation in terms of inference time, when applying compiler optimizations.
@@ I Introduction
Overall, we find that conversions between DL frameworks significantly impacts output labels of the DNN models by up to 100%. We identify the source of the label-impacting error – small amounts of noise introduced in the weights by the framework conversion tools. The weight differences, although small, may be caused by floating-point rounding errors which can cause label changes when accumulated across the layers. On the other hand, we found that varying hardware accelerators and compiler optimizations do not affect model output but can lead to a non-negligible performance degradation with respect to inference time under specific scenarios. We observed up to 81% unexpected performance degradation in model inference times when applying certain compiler optimizations.
@@ VI-A1 RQ1: Label Sensitivity to DL Framework Conversions
The results are presented in Figure 6, showing the degree of dissimilarity between source and target models. As can be seen from the empty gray boxes, the conversion tool crashes in 11 out of the 36 conversions across the three DNN models, indicating that the conversion process failed. This happened due to compatibility issues between the conversion tool and a given model architecture, or the source or target DL framework. For instance, we could not convert any Keras models to TensorFlow, due to the tf2onnx tool being unable to handle some tensor element values. Additionally, we observe further 15 cases where the conversion succeed without crashing, but the target model gave different labels from the source model, with 11 of these cases having a significant discrepancy (over 35%). In particular, we observe a 100% discrepancy in the output labels when converting the ResNet101 model from TFLite to Keras.
@@
Finally, the conversion of TF models to Keras gives varying results across models, with MobileNetV2 having 49% dissimilarity, ResNet101 having 35%, and InceptionV3 giving a model crash. This points to weaknesses in the conversion tool with certain DNN model architectures.
@@
We use the Analysis part of DeltaNN to explore in greater detail the cause of discrepancies across DNN model conversions. To illustrate the analysis in depth, we select one of the models and conversions that results in a discrepancy, InceptionV3 with TF as the source DL framework converted to target TFLite. The discrepancies observed across source-target is 4%, and in Figure 7 we show the class breakdown of the images which demonstrated differences, sorted by what proportion of that class showed discrepancies. We highlight a subset of the class labels on the x-axis. We observe that some classes are impacted more than others, with some classes such as “walker hound” disagreeing on 100% of the images. However, this graph is not indicating test set accuracy, instead it is about agreement between the source and converted target models, which under ideal circumstances we would expect to have equal agreement in all cases.
@@ VIII Conclusion
We introduced the DeltaNN Differential testing framework to explore the impact of computational environment parameters on image recognition models. In particular, we study the effect of converting between popular deep learning frameworks (TensorFlow, Keras, TFLite, PyTorch), compiler optimization settings, and hardware devices on the output labels of three widely used image recognition models. We also monitor the impact of these parameters on model inference time. Overall, we find that conversions between DL frameworks significantly impact output labels of the DNN models by up to 100%. Our framework also provides analysis capabilities for label discrepancies stemming from framework conversions.